Ticket #4 (closed defect: fixed)

Opened 13 years ago

Last modified 13 years ago

Andrew's old A-service still visible on a number of datasets

Reported by: lawrence Owned by: selatham
Priority: desirable Milestone: Replace Metadata Gateway
Component: discovery Version:
Keywords: Cc:

Description

We need to clean up the discovery records that point to the old e-science service that no longer works, e.g.  http://dmgdev1.esc.rl.ac.uk/cgi-bin/ndgDataAccess?datasetSource=dmgdev1.esc.rl.ac.uk&datasetID=acsoe_badc_eae96_macehead_mt1_1

Change History

comment:1 Changed 13 years ago by lawrence

  • Priority changed from minor to desirable

comment:2 Changed 13 years ago by selatham

  • Status changed from new to closed
  • Resolution set to fixed

Simply deleted the 'methyl chloroform' and 'meteorology' dataset DIFs from the live service eXist db. These are test sets of MOLES objects & D metadata created at BADC. They should remain in the Development eXist database for testing purposes. The links to A services can be changed as required for testing further developments.

Note: See TracTickets for help on using tickets.